• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Jen

    Log entry time 15:46:32 on April 4,2010

    Entry number 316184

    keyword=04/04/10 Day Shift Summary

    05:50 AM - 08:15 AM L-arm Q1 quench problem
    08:25 AM Bogdan Wojtsekhowski called Mark to ask if we need to give a
    call for help. Mark said he will call Jack to make a
    decision.
    09:07 AM Jon finished his A_T scan study. Rupesh is on the way to
    JLab. Compton people are waiting for starting the
    commssioning.
    09:23 AM move the target to the empty; turn the septum off; Rupesh is
    working on BCM/Lumi quality study with Parity DAQ. Jon stays
    around to help Rupesh.
    09:33 AM ask MCC to send 50 uA beam
    10:24 AM Rupesh requests for the adjustment of Lumi HV; 150k Lumi
    signal on thick Carbon target
    10:57 AM ask MCC to turn the beam off; then, turn the septum on.
    Rupesh couldn't adjust HV for det#1 (Upper Quartz), det#2
    (Lower Quartz), and A_T Quartz. Rupesh called Luis. Jon is
    sleeping.
    11:05 AM ask MCC to turn the beam on; keep the same target, the same
    raster size
    11:56 AM adjust Beam Lumi HV for 150k HV signals on thick Carbon
    target with 50uA (Septum on)
    12:00 PM - Rupesh is working on beam study
    02:46 PM ask MCC to turn the beam off
    02:48 PM ask MCC to turn the beam on (50uA)
    03:22 PM ask MCC to turn the beam off; move the target to empty; turn
    the septum on
    03:46 PM hand over to the next shift leader